OpenStudy (anonymous):

A softball is thrown up into the air. The ball's height in feet is modeled by the equation h(t) = -16t2 + 32t + 5. h = height t = time in seconds Find the maximum height of the ball. a) 13.9 feet b) 8.3 feet c) 12.2 feet d) 7.6 feet

OpenStudy (anonymous):

should be 21 feet, try plot it on ur calculator. All ur options given seem to be incorrect.

OpenStudy (anonymous):

Note that the height function is a quadratic. The vertex of the parabolic graph for such a function is at x=-b/2a=1. The maximum value for the function occurs at this point, i.e., h(1)=-16+32+5=21. As far as I can tell, the correct answer is not listed.